If x represents the number of voys in a family of 7 children, find the probabbility and distribution

1
Expert's answer
2022-05-03T16:24:05-0400

"X\sim Bin(n, p)"

"n=7, p=q=0.5."


"P(X=k)=\dbinom{7}{k}(0.5)^k(0.5)^{7-k}=\dfrac{\dbinom{7}{k}}{128}"


"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 x & p(x)\\ \hline
 0 & 1/128\\ \hdashline
 1 & 7/128\\ \hdashline
 2 & 21/128\\ \hdashline
 3 & 35/128\\ \hdashline
 4 & 35/128\\ \hdashline
 5 & 21/128\\ \hdashline
 6 & 7/128\\ \hdashline
 7 & 1/128\\ \hdashline
\end{array}"